10. If a pump discharges 90 GPM of water whose specific weight is 61.5 lb f /ft 3 (g = 31.95 ft/s 2 ), find a) mass

flow rate (lb m /min), b) time to fill a spherical tank 8 ft in diameter.


1
Expert's answer
2021-09-10T00:04:03-0400

a)

There are "7.48 ft^3" in a gallon

Therefore, the volume flow rate can be expressed as "\frac{75 gallon/min}{7.48 gallon /ft^3}= 10.03 ft^3/min"

We can multiply that by the density of 62.4 lbm/ft^3

Mass flow rate = 626 LBM/min


b)

"Time = \frac{Volume}{Flowrate}\\
=\frac{\pi}{4}*\frac{L}{10.03}ft^3/min\\
= 94.0 minutes"